OpenStudy (buttermequeasy):

John is going to rent a tent to go camping. The cost to rent a camping tent from Woodland Outfitters is a flat rate of $12.00 plus $2.00 for each night the tent is rented. Part a. Let x represent the number of nights John keeps the tent and write a cost function, c(x), for the cost to rent a tent from Woodland Outfitters. Part b. How much will John owe if he uses the tent for three nights?

OpenStudy (karatechopper):

well hello there such a beautiful day today right? NOW BECOME A V OR ELSE....

OpenStudy (jamesj):

So, part 1. If he has it one night, then the cost c(1) = 12 + 2 two nights: c(2) = 12 + 2 + 2 = 12 + 2(2) hence for x nights, c(x) = 12 + 2x

OpenStudy (jamesj):

Now part 2. The cost for three nights is c(3), what's the value of c(3)?

OpenStudy (jamesj):

butter, you still there?

OpenStudy (buttermequeasy):

Yes Sorry

OpenStudy (jamesj):

Does the function c(x) make sense?

OpenStudy (buttermequeasy):

Yes.